A PAF helicopter is flying with a horizontal speed of 8.00 m/s distributing relief goods to lahar stricken families in Pampanga. A family stranded on the roof of their house is about 6.00 m below the chopper. With what velocity does the package hit the roof? *




1
Expert's answer
2021-11-19T11:10:11-0500

The velocity is


"v=\sqrt{v_x^2+v_y^2}=\sqrt{8^2+2gh}=13.5\text{ m/s},\\\space\\
\theta=\arctan\frac{\sqrt{2gh}}{v_x}=54°"

below horizontal.